How would you factor x^2 + 9x + 18 ? What are the steps?
Found 2 solutions by Fombitz, Timnewman:
Answer by Fombitz(32388)

Look for factors of 18 that sum to 9.
(1,18)
(2,9)
(3,6)

Answer by Timnewman(323)

hi dear,
To factorize x²+9x+18,
multiply the last term with the fist term,i.e x²*18=18x².
Find two number such that their product is 18x² and their sum gives 9x (i.e the second term);
Replace the the middle term(9x) with the sum of the the two numbers found.then group the expression by the use of bracket nd then fctorize.
HERE IS THE WORKING.:
6x*3x=18x²
6x+3x=9x
replace the middle term with 6x+3x in the given expression
(x²+6x)+(3x+18)
factor each bracket
x(x+6)+3(x+6)
then the factors
(x+3)(x+6)
--TIM